OpenStudy (harleyquinn):

4 problems: given f(x)=2xt3 and g(x)=x(sq)-5x evaluate f(4) evaluate f(6m) evaluate g(-7) evaluate g(t+8)

OpenStudy (anonymous):

F(4)=2(4)+3=11 F(6m) = 2(6m)+3=12m+3

OpenStudy (anonymous):

the above is correct. same goes for g(x) substitute the values then solve.

OpenStudy (anonymous):

g(x)=x(sq)-5x evaluate g(-7) g(-7) = (-7)^2 - 5(-7) g(-7) = 49 + 35 evaluate g(t+8) g(t + 8) = (t+8)^2 - 5(t + 8) y = t^2 + 16t + 64 - 5t - 40

OpenStudy (harleyquinn):

Thanks ^-^